XRCF 2ND FLOOR SHIFT REPORT Starting Date: 16 January 1997 Shift: A File Ref: /home/xrcfdude/shiftreports/970116.A RS/RE: EMK Operator: JFM Software: JFM Shift Scientist: INE Project Scientist: JK ==================== Plan: ---- Grating 1D scans. Measurement Summary: ------------------- 109056 D-HXF-P1-19.001a PSF/1D, HEG 0 order 109057 D-HXF-P1-19.002a PSF/1D, HEG +1 order 109058 D-HXF-P1-19.003 PSF/1D, HEG +2 order 109059 D-HXF-P1-19.004 PSF/1D, MEG 0 order, beamcen 109060 D-HXF-P1-19.004 PSF/1D, MEG 0 order, beamcen in Y 109061 D-HXF-P1-19.004 PSF/1D, MEG 0 order 109062 D-HXF-P1-19.005 PSF/1D, MEG +1 order 109063 D-HXF-P1-19.006 PSF/1D, MEG +3 order 109064 D-HXF-P1-19.007 PSF/1D, HEG 0 order 109065 D-HXF-P1-19.008 PSF/1D, HEG +1 order 109066 D-HXF-P1-19.009 PSF/1D, HEG +2 order 109068 D-HXF-P1-19.010 PSF/1D, MEG 0 order 109069 D-HXF-P1-19.011 PSF/1D, MEG +1 order Notes: ----- We re-ran D-HXF-P1-19.001 and 19.002 as 19.001a and 19.002a after some manual massaging of the procedures to (a) remove a focusmove that would have lost the results of the preceding beamcenter, (b) modifying the make-line to use the starting position from the beamcenter, and (c) modifying the beamcenter step size to be appropriate for the long axis of the slit apertures. Similar changes were made for 19.003-19.006. JFM revised the ccl procedure to behave correctly, and we re-installed 19.007-19.018 and also D-LXF-P1-17.001-17.009. We skipped the beamcen's in 19.001a, 19.002a, and 19.003, using instead the values previously measured from 19.002 and the nominal offsets. Likewise, the beamcen's in 19.005 and 19.006 were skipped, using the nominal offsets from the beamcen in 19.004. The beamcen's in 19.007-19.012 were skipped because the aperture is large enough to not require beamcentering. For D-HXF-P1-19.004 the (2D) beamcen would not converge in Y because the Y position of the peak depends on the Z position in the slanted slits. These should be run as pairs of 1D beamcens. Run 109067 (D-HXF-P1-19.010) was aborted when it was determined that the primey stage was not in the correct position. The test was restarted as 109068. Problems: -------- Shell 3 north shutter failed to open twice (1535Z and 1853Z) in addition to once near the end of the previous shift. Closing all shutters and reexecuting the open command resolved the problem in each case. We had several cases of bad IRIG timing at the BND5 PC. At 1420Z we found that the PC was not receiving an IRIG signal from the facility. Coincidentally the BND500 gas system experienced an anomaly that resulted in a valve closing and ~20 torr pressure drop. The IRIG signal returned a few minutes later. We had two cases of child processes not returning; they had to be killed manually. XRCF 2ND FLOOR SHIFT REPORT Starting Date: 16 January 1997 Shift: B File Ref: /home/xrcfdude/shiftreports/970116.B RS/RE: W. A. Podgorski Operator: T. Norton Software: J. Moran Shift Scientist: G. Zirnstein Project Scientist: M. Weisskopf ==================== Plan: ---- Continue measurements on EIPS Mg-K Measurement Summary: -------------------- D-HXF-P1-19.012 PSF/1D MEG 3rd order 109070 listscan 3rd order D-HXF-P1-19.016 PSF/1D HETG, 0th order 109071 beamcen 0th order 109072 listscan at 0th order, process hung, killed child process D-HXF-P1-19.017 PSF/1D HETG, 1st order move to 0 order done AFTER ei move to 1st order, resulting in bad listscan at 0th order, out of focus!! 109073 listscan, process hung, killed child process D-HXF-P1-19.018 PSF/1D METG, 3rd order 109074 beamcen 3rd order(did not converge) focusmove to 0th order MEG absolute move to 0th order meg from 19.016 incremental move to 3rd order 109075 listscan 3rd order D-HXF-P1-19.017a Repeat PSF/1D at 1st order focusmove to meg 0th order absolute move to 0th order center from 19.016 incremental move to 1st order 109076 listscan 1st order D-HXF-P1-19.013 PSF/1D HETG (4 & 6 shells) 0th order 109077 beamcen 0th order 109078 listscan 0th order D-HXF-P1-19.014 PSF/1D HETG 1st order focusmove fpc_x2 AD heg move absolute to 0th order (from 19.013) move incremental to 1st order 109079 listscan at 1st order D-HXF-P1-19.015 PSF/1D HETG, 2nd order focusmove fpc_x2 AD heg move absolute to 0th order (from 19.013) move incremental to 2nd order 109080 listscan 2nd order D-HXF-P1-17.001 PSF/1D LETG, 0th order shutter 3N stuck, reissued command to open focusmove fpc_x2 AA leg 109081 beamcen 0th order 109082 listscan 0th order D-HXF-P1-17.002 PSF/1D LETG, 1st order focusmove fpc_x2 AA leg absolute move to 0th order from 17.001 incremental move to 1st order 109083 listscan 1st order D-HXF-P1-17.003 PSF/1D LETG, 5th order focusmove fpc_x2 AA leg absolute move to 0th order from 17.001 incremental move to 5th order 109084 listscan 5th order D-HXF-P1-17.004 PSF/1D LETG, 0th order, all shutters 109085 listscan 0th order Notes: An index of fpc_x2 aperture updates was created and put in ----- shift report book, along with diffs on tables from update to update Problems: Shutter 3N stuck once on thes 17.001 -------- SAO-OP1 crashed and had to be power cycled to restart There were two occurences of hung listscans that were cleared by killing the child process. XRCF 2ND FLOOR SHIFT REPORT Starting Date: 16 January 1997 Shift: C File Ref: /home/xrcfdude/shiftreports/970116.C RS/RE: B. Podgorski/D. Schwartz Operator: N. Mistry Software: N. Mistry Shift Scientist: B. Wargelin Project Scientist: L. Van Speybroeck ==================== Measurement Summary: ------------------- 109085 D-LXF-P1-17.004 LETG 0th order 80x500v slit scan in y 109086 D-LXF-P1-17.005 LETG 1st order 80x500v slit scan in y 109087 D-LXF-P1-17.006 LETG 5th order 80x500v slit scan in y 109088 D-LXF-P1-17.007 LETG 0th order 10x200h slit scan in z 109089 D-LXF-P1-17.008 LETG 1st order 10x200h slit scan in z 109090 D-LXF-P1-17.009 LETG 5th order 10x200h slit scan in z